What must be the magnitude of the displacement of the automobile (from rest) after 4 seconds of travel if its acceleration as function time is a(t) = 2t m/s^2 along positive x-axis?


1
Expert's answer
2021-09-20T09:59:37-0400

Let's assume, the car starts at "x_0=0" with zero initial speed "v_0=0". Then the velocity of the car is given as follows:


"v(t) = v_0 + \int_0^ta(t)dt = \int_0^t2tdt = t^2\space m/s"

The displacement is:


"s(t)=x(t)-x_0 = x(t) = \int_0^tv(t)dt = \int_0^tt^2dt = \dfrac{t^3}{3}\space m"

Substituting "t = 4s", obtain:


"s(4) = \dfrac{4^3}{3}= \dfrac{64}{3} m\approx 21.3m"

Answer. Approximately 21.3 m.